OpenStudy (anonymous):

Medal~ (2^8 ⋅ 3^−5 ⋅ 6^0)^−2

OpenStudy (anonymous):

Exact Form: 59049/3839537296 Decimal Form: 1.53791969⋅10−^5

OpenStudy (anonymous):

Thank you, do you think you can show me how you did it? ouo

OpenStudy (anonymous):

I understand how to do two, but not three

OpenStudy (anonymous):

@MaydayPaRAYde

OpenStudy (anonymous):

Aha Sure ^-^ @CursedAura

OpenStudy (anonymous):

Given f(x) = 3x + 2 and g(x) = 4 – 5x, find (f + g)(x), (f – g)(x), (f×g)(x), and (f / g)(x). To find the answers, all I have to do is apply the operations (plus, minus, times, and divide) that they tell me to, in the order that they tell me to. (f + g)(x) = f(x) + g(x) = [3x + 2] + [4 – 5x] = 3x – 5x + 2 + 4 = –2x + 6 (f – g)(x) = f(x) – g(x) = [3x + 2] – [4 – 5x] = 3x + 5x + 2 – 4 = 8x – 2 (f×g)(x) = [f(x)][g(x)] = (3x + 2)(4 – 5x) = 12x + 8 – 15x2 – 10x = –15x2 + 2x + 8 (f/g)(x) = [f(x)]/[g(x)] = [3x + 2]/[4 - 5x] Given f(x) = 2x, g(x) = x + 4, and h(x) = 5 – x3, find (f + g)(2), (h – g)(2), (f × h)(2), and (h / g)(2). To find the answers, I can either work symbolically (like in the previous example) and then evaluate, or I can find the values of the functions at x = 2 and then work from there. It's probably simpler in this case to evaluate first, so: f(2) = 2(2) = 4 g(2) = (2) + 4 = 6 h(2) = 5 – (2)3 = 5 – 8 = –3 Now I can evaluate the listed expressions: (f + g)(2) = f(2) + g(2) = 4 + 6 = 10 (h – g)(2) = h(2) – g(2) = –3 – 6 = –9 (f × h)(2) = f(2) × h(2) = (4)(–3) = –12 (h / g)(2) = h(2) ÷ g(2) = –3 ÷ 6 = –0.5 If you work symbolically first, and plug in the x-value only at the end, you'll still get the same results. Either way will work. Evaluating first is usually easier, but the choice is up to you.

OpenStudy (anonymous):

This Video explains it better love, I am not good at explaining https://www.khanacademy.org/math/algebra/solving-linear-equations-and-inequalities/old-school-equations/v/order-of-operations-example

OpenStudy (anonymous):

Here's another one incase that one didn't get you through! https://www.khanacademy.org/math/algebra/introduction-to-algebra/variable-and-expressions/v/evaluating-exponential-expressions-2